OpenStudy (anonymous):

find five consecutive multiples of three whose sum is 210?

OpenStudy (callisto):

let x be the smallest multiple. Then the others terms are x+3, x+6, x+9, and x+12 Sum = 210 x + (x+3) + (x+6) + (x+9) +(x+12) =210 5x + 30 = 210 Can you solve x?

OpenStudy (anonymous):

Ok thank you so much:D

OpenStudy (callisto):

Hmm.. perhaps it's better to say let x be the smallest number required .... :S

OpenStudy (callisto):

welcome!~